# WiMANS: A Benchmark Dataset for WiFi-based Multi-user Activity Sensing

### 24 Jan 2024 v1

@@ 3.2 Dataset Statistics
Figure 5: Statistics of WiMANS regarding the distributions of identities, locations, activities, and numbers of users per sample ($N_{u}$). In each distribution, all categories have almost equivalent proportions to construct a relatively balanced dataset.
Data Distribution. WiMANS incorporates varying numbers of users in 11286 samples. 594 samples (5.3%) contain 0 user; 3564 samples (31.6%) contain 1 user; 7128 (63.2%) samples contain multiple users. Among the multi-user samples, there are 1782 (15.8%) samples for 2 users, 1782 (15.8%) samples for 3 users, 1782 (15.8%) samples for 4 users, and 1782 (15.8%) samples for 5 users. Figure 5 presents the statistics of WiMANS. There are 4752 samples for each user, 5778$\sim$5886 samples for each location, and 3168 samples for each activity, empowering models to learn representative features for different tasks. Note that the sum of identities/locations/activities is not equal to the total sample number (11286), since each multi-user sample involves more than one identity, location, and activity. Packet Loss. Ideally, each CSI sample consists of 3000 time steps, as mentioned in Section 3.1. However, packet loss inevitably exists in wireless communication, leading to missing time steps in CSI samples. The average packet loss rates are 4.52% and 2.31% for 2.4 GHz band and 5 GHz band, respectively. 2.4 GHz band suffers from more severe packet loss because it is more commonly used and more crowded than 5 GHz band, affected by more environmental noise [55, 51]. Thus, most existing works [60, 64, 2] apply 5 GHz band for WiFi-based human sensing. However, compared to 5 GHz signals, 2.4 GHz signals can theoretically cover a larger area using longer wavelength to better penetrate and/or bypass obstacles. This inspires us to study dual-band augmented sensing in the future, since WiMANS has gathered CSI of both 2.4 GHz and 5 GHz WiFi bands.
@@ 4 Experiments
2.4 GHz
ST-RF [60] 66.31$\pm$1.33 62.89$\pm$0.36 56.82$\pm$0.14 78.99$\pm$1.32 65.69$\pm$0.55 56.75$\pm$0.09 71.01$\pm$2.20 61.46$\pm$0.54 56.66$\pm$0.09
MLP [58] 65.70$\pm$0.91 60.46$\pm$0.47 57.56$\pm$0.13 75.86$\pm$0.62 64.21$\pm$0.47 57.80$\pm$0.16 75.33$\pm$0.59 58.82$\pm$0.26 57.03$\pm$0.12
LSTM [60] 80.74$\pm$1.43 66.91$\pm$1.09 59.16$\pm$0.42 89.52$\pm$1.11 71.72$\pm$0.88 58.95$\pm$0.33 86.68$\pm$0.99 67.84$\pm$0.65 57.72$\pm$0.54
CNN-1D [49] 84.80$\pm$0.80 73.23$\pm$0.22 59.55$\pm$0.24 92.89$\pm$0.26 77.21$\pm$0.19 58.47$\pm$0.11 88.14$\pm$0.46 74.51$\pm$0.31 58.02$\pm$0.17
CNN-2D [34] 89.97$\pm$0.66 76.98$\pm$0.37 59.54$\pm$0.15 96.39$\pm$0.81 82.10$\pm$0.40 58.99$\pm$0.12 91.11$\pm$0.82 79.08$\pm$0.52 58.17$\pm$0.26
CLSTM [33] 90.16$\pm$0.70 72.97$\pm$0.68 61.79$\pm$0.55 94.56$\pm$0.49 77.42$\pm$0.75 60.87$\pm$0.39 92.76$\pm$0.27 71.80$\pm$0.95 60.97$\pm$0.40
ABLSTM [6] 88.81$\pm$1.36 76.57$\pm$0.42 61.57$\pm$0.50 94.80$\pm$0.55 77.71$\pm$0.61 60.86$\pm$0.27 90.90$\pm$0.54 75.98$\pm$0.55 59.66$\pm$0.15
THAT [28] 90.69$\pm$0.51 78.78$\pm$0.74 61.04$\pm$0.32 94.77$\pm$0.59 82.21$\pm$0.84 60.14$\pm$0.28 93.87$\pm$0.88 80.83$\pm$1.11 59.71$\pm$0.29
5 GHz
ST-RF [60] 89.36$\pm$1.08 62.72$\pm$0.57 57.31$\pm$0.08 95.09$\pm$0.83 67.41$\pm$0.65 57.58$\pm$0.13 81.86$\pm$1.26 61.93$\pm$0.56 57.21$\pm$0.21
MLP [58] 98.57$\pm$0.13 72.67$\pm$0.82 58.58$\pm$0.14 99.52$\pm$0.11 79.10$\pm$0.42 57.38$\pm$0.14 95.09$\pm$0.58 74.45$\pm$0.87 59.18$\pm$0.23
LSTM [60] 98.89$\pm$0.16 76.13$\pm$0.68 60.58$\pm$0.28 99.73$\pm$0.17 81.30$\pm$0.80 59.49$\pm$0.23 95.84$\pm$0.50 77.43$\pm$0.37 59.84$\pm$0.36
CNN-1D [49] 99.47$\pm$0.21 82.41$\pm$0.34 60.56$\pm$0.22 99.66$\pm$0.12 87.13$\pm$0.32 59.03$\pm$0.16 95.94$\pm$0.36 83.95$\pm$0.58 60.16$\pm$0.14
CNN-2D [34] 99.10$\pm$0.21 82.71$\pm$0.41 59.78$\pm$0.30 99.79$\pm$0.11 88.20$\pm$0.37 58.58$\pm$0.29 95.31$\pm$0.36 84.36$\pm$0.40 59.55$\pm$0.28
CLSTM [33] 99.71$\pm$0.19 81.93$\pm$0.50 64.24$\pm$0.55 99.76$\pm$0.08 88.07$\pm$0.55 62.06$\pm$0.36 97.64$\pm$0.25 84.35$\pm$0.64 64.82$\pm$0.38
ABLSTM [6] 99.55$\pm$0.29 82.53$\pm$0.67 61.40$\pm$0.22 99.76$\pm$0.08 87.44$\pm$0.25 60.43$\pm$0.16 96.74$\pm$0.38 83.75$\pm$0.53 61.32$\pm$0.15
THAT [28] 99.23$\pm$0.19 83.09$\pm$0.67 61.75$\pm$0.29 99.99$\pm$0.00 88.06$\pm$0.59 61.15$\pm$0.37 97.08$\pm$0.29 83.66$\pm$1.05 62.13$\pm$0.41
2.4 / 5 GHz
ST-RF [60] 77.89$\pm$1.05 62.46$\pm$0.39 57.33$\pm$0.09 87.72$\pm$0.91 66.69$\pm$0.34 57.39$\pm$0.07 75.99$\pm$1.06 62.10$\pm$0.28 57.21$\pm$0.08
MLP [58] 74.94$\pm$0.48 66.01$\pm$0.32 57.94$\pm$0.13 85.90$\pm$0.78 69.23$\pm$0.28 57.62$\pm$0.07 80.12$\pm$0.53 66.17$\pm$0.48 58.44$\pm$0.08
LSTM [60] 86.93$\pm$0.90 69.53$\pm$0.79 59.58$\pm$0.24 93.94$\pm$0.60 75.19$\pm$0.46 59.66$\pm$0.31 90.07$\pm$0.95 70.37$\pm$0.27 58.90$\pm$0.25
CNN-1D [49] 92.64$\pm$0.42 78.25$\pm$0.25 60.72$\pm$0.14 96.64$\pm$0.21 82.75$\pm$0.25 59.52$\pm$0.17 93.49$\pm$0.21 78.51$\pm$0.21 59.68$\pm$0.12
CNN-2D [34] 94.61$\pm$0.50 79.21$\pm$0.40 59.16$\pm$0.29 98.01$\pm$0.35 83.58$\pm$0.36 58.50$\pm$0.28 94.45$\pm$0.41 80.07$\pm$0.51 58.26$\pm$0.21
CLSTM [33] 93.05$\pm$0.46 79.90$\pm$0.48 63.95$\pm$0.42 97.44$\pm$0.28 83.12$\pm$0.33 62.38$\pm$0.28 95.29$\pm$0.38 78.23$\pm$0.60 63.46$\pm$0.36
ABLSTM [6] 92.35$\pm$0.39 79.17$\pm$0.38 62.14$\pm$0.16 97.84$\pm$0.17 82.81$\pm$0.42 60.92$\pm$0.17 95.01$\pm$0.34 78.88$\pm$0.54 61.43$\pm$0.29
THAT [28] 94.29$\pm$0.51 80.42$\pm$0.56 61.75$\pm$0.19 97.73$\pm$0.33 83.25$\pm$0.45 61.22$\pm$0.22 95.80$\pm$0.40 79.97$\pm$0.82 61.95$\pm$0.33
Table 2: Recognition performance of WiFi-based models on WiMANS in terms of accuracy (%).
This section utilizes WiMANS to benchmark the multi-user sensing performance of WiFi-based and video-based models, with respect to human identification, localization, and activity recognition (HAR). We also compare these models regarding model complexity and time efficiency.
4.1 Experimental Setup
Baselines. We evaluate 8 WiFi-based models on WiMANS, including Random Forest based on Short-time Fourier Transform (ST-RF) [60], MLP [58], LSTM [60], CNN-1D [49], CNN-2D [34], CLSTM [33], ABLSTM [6], and THAT [28]. Particularly, CLSTM [33], CNN-1D [49], and THAT [28] have achieved state-of-the-art performance in WiFi-based human identification, localization, and activity recognition, respectively. Meanwhile, we apply 6 state-of-the-art video classification models for comparison, including ResNet [43], S3D [53], MViT-v1 [12], MViT-v2 [29], Swin-T [31], and Swin-S [31]. Evaluation Metrics. We employ accuracy to measure the recognition performance of multi-user sensing, following the previous works [60, 6, 28]. Note that each sample in WiMANS includes multiple labels (identities/locations/activities), and thus we measure the accuracy of recognizing labels rather than samples. To measure model complexity and time efficiency, we adopt the number of parameters, floating point operations (FLOPs), and recognition throughput as metrics. Implementation Details. We analyze model performance in different environments and accordingly split WiMANS into 3 subsets. Each subset is randomly split into a training set (80%) and a test set (20%). To recognize multiple labels, we connect each model to a linear layer followed by a sigmoid function for multi-label classification. For WiFi-based models, we leverage publicly accessible implementations as much as possible and initialize them with Xavier [16]. For video-based models, we utilize the implementations provided by PyTorch and initialize them with the weights pre-trained by Kinetics-400 [4]. Inputs of all models are resized following the original papers (e.g., 90$\times$3$\times$112$\times$112 for ResNet [43]). We repeat each experiment 10 times with random seeds and report the means and standard deviations of results. More details are provided in Appendix C.
4.2 Comparison Results
@@
ResNet [43] 99.04$\pm$0.93 95.93$\pm$1.70 72.32$\pm$0.86 99.20$\pm$0.94 95.60$\pm$5.14 73.62$\pm$1.13 96.08$\pm$1.39 96.18$\pm$5.38 76.53$\pm$4.74
S3D [53] 99.99$\pm$0.00 99.54$\pm$0.50 94.00$\pm$1.35 99.85$\pm$0.04 99.92$\pm$0.02 98.01$\pm$0.12 99.73$\pm$0.32 99.95$\pm$0.01 98.51$\pm$0.15
MViT-v1 [12] 99.97$\pm$0.05 99.26$\pm$0.26 93.70$\pm$0.44 99.73$\pm$0.12 99.56$\pm$0.05 90.72$\pm$0.11 99.72$\pm$0.04 99.79$\pm$0.01 94.03$\pm$0.10
MViT-v2 [29] 99.83$\pm$0.06 99.71$\pm$0.05 95.12$\pm$0.11 99.76$\pm$0.13 99.77$\pm$0.04 92.29$\pm$0.13 99.80$\pm$0.12 99.85$\pm$0.01 95.18$\pm$0.11
Swin-T [31] 99.76$\pm$0.05 99.59$\pm$0.05 96.55$\pm$0.12 99.97$\pm$0.05 99.73$\pm$0.07 89.06$\pm$0.14 99.85$\pm$0.04 99.81$\pm$0.01 94.09$\pm$0.13
Swin-S [31] 99.95$\pm$0.07 99.59$\pm$0.06 95.84$\pm$0.11 99.50$\pm$0.08 99.44$\pm$0.05 90.51$\pm$0.11 99.83$\pm$0.12 99.75$\pm$0.02 96.05$\pm$0.11
Table 3: Recognition performance of video-based models on WiMANS in terms of accuracy (%).
@@
Table 4: Model complexity and time efficiency on WiMANS. (“Recs/s”: recognitions per second.)
WiFi-based Multi-user Sensing. Table 2 presents the multi-user sensing performance of WiFi-based models. (1) Using the 2.4 GHz WiFi band, THAT and CNN-2D are the best performing models for identification, while CNN-2D outperforms other models in localization, and CLSTM achieves the best performance in HAR. (2) With the 5 GHz WiFi band, CLSTM and THAT show the highest accuracy in recognizing multiple identities, while THAT and CNN-2D demonstrate their superiority in localization, and CLSTM obtains better results than other models in HAR. (3) Exploiting both 2.4 and 5 GHz WiFi bands, CNN-2D and THAT have the best results in identifying and localizing multiple users, while CLSTM yields the highest accuracy for HAR. Generally, we can observe that the state-of-the-art models show better performance on identification (90.69$\sim$99.99%) than localization (78.78$\sim$88.20%), while having even lower performance on HAR (60.87$\sim$64.82%). This is because HAR is more fine-grained than the other two tasks, causing higher difficulties to learn representative features. In contrast, multi-user identification is less fine-grained than localization and thereby results in better performance. Comparing different WiFi bands, using 5 GHz results in higher accuracy than 2.4 GHz, since the 2.4 GHz WiFi band suffers from more noise than 5 GHz, as mentioned in Section 3.2. Meanwhile, simply using the samples of two WiFi bands cannot produce models of better performance. These inspire us to study dual-band augmented multi-user sensing in the future. Numbers of Users. Figure 6 analyzes the impacts of user numbers on the performance of THAT. As expected, sensing more users simultaneously results in lower performance. For example, when the maximum user number increases from 1 to 5 in the classroom, the results of identification, localization, and activity recognition decrease by 5.7%, 15.4%, and 25.74%, respectively. These results illustrate the usefulness of WiMANS and the challenges of WiFi-based multi-user sensing. Video-based Multi-user Sensing. We provide the performance of video-based models in Table 3 for comparison. Overall, video-based models achieve better performance than WiFi-based models in three sensing tasks. Such results are expected since video-based analysis has been well studied in computer vision, which also demonstrate the quality of WiMANS, where users have performed distinguishing activities. Despite the remarkable performance of video-based models, they have much higher complexity and lower efficiency than WiFi-based models, as discussed below. Model Complexity and Time Efficiency. Table 4 highlights the superiority of WiFi-based models compared with video-based models regarding model complexity and time efficiency. Except for MLP which is rarely used in practice due to excessive parameters, the maximum number of parameters (5.391 M for CLSTM) in WiFi-based models is 1.55$\times$ less than the minimum number of parameters (8.342 M for S3D) in video-based models. Similarly, the maximum FLOPs (3.208 G for ABLSTM) in WiFi-based models is 5.51$\times$ lower than the minimum FLOPs (17.670 G for ResNet) in video-based models. Moreover, the throughput of THAT is 16.33$\sim$17.02$\times$ higher than the throughput of S3D. We further will discuss the training time and testing time of models in Appendix C.

### 12 Mar 2024 v2

@@ 3.2 Dataset Statistics
Figure 4: Statistics of WiMANS regarding the distributions of user identities, locations, activities, and numbers of users per sample. In each distribution, all categories have almost equivalent proportions to construct WiMANS as a relatively balanced dataset.
Data Distribution. WiMANS incorporates varying numbers of users in 11286 samples. 594 samples (5.3%) contain 0 user; 3564 samples (31.6%) contain 1 user; 7128 (63.2%) samples contain multiple users. Among the multi-user samples, there are 1782 (15.8%) samples for 2 users, 1782 (15.8%) samples for 3 users, 1782 (15.8%) samples for 4 users, and 1782 (15.8%) samples for 5 users. Fig. 4 presents the statistics of WiMANS. There are 4752 samples for each user, 5778$\sim$5886 samples for each location, and 3168 samples for each activity, empowering models to learn representative features for different sensing tasks. Note that the sum of identities/locations/activities is not equal to the total sample number (11286), since each multi-user sample involves more than one identity, location, and activity. Packet Loss. Ideally, each CSI sample consists of 3000 time steps, as mentioned in Section 3.1. However, packet loss inevitably exists in wireless communication, leading to missing time steps in CSI samples. The average packet loss rates are 4.52% and 2.31% for the 2.4 GHz band and the 5 GHz band, respectively. On average, 2.4 GHz samples have 2864.39 time steps, while 5 GHz samples have 2930.72 time steps. 2.4 GHz band suffers from more severe packet loss because it is more commonly used and more crowded than 5 GHz band, affected by more environmental noise [71, 62]. Thus, most existing works [73, 77, 3] only apply the 5 GHz band for WiFi-based human sensing. However, compared to 5 GHz signals, 2.4 GHz signals can theoretically cover a larger area using longer wavelength to better penetrate and/or bypass obstacles. This inspires us to study dual-band augmented human sensing in the future, since WiMANS has gathered CSI of both 2.4 GHz and 5 GHz WiFi bands.
@@ 4 Experiments
This section utilizes WiMANS to benchmark the multi-user sensing performance of state-of-the-art WiFi-based and video-based models, with respect to human identification, localization, and activity recognition (HAR). We also compare these models in terms of model complexity and time efficiency.
4.1 Baselines
WiFi-based models. We evaluate 8 WiFi-based models on WiMANS, including Random Forest based on Short-time Fourier Transform (ST-RF) [73], MLP [67], LSTM [73], CNN-1D [58], CNN-2D [42], CLSTM [41], ABLSTM [7], and THAT [34]. Particularly, CLSTM [41] is a hybrid of CNNs and LSTM, outperforming other models in WiFi-based identification. For WiFi-based localization, CNN-1D [58] has demonstrated the best performance by learning human features along the temporal dimension. THAT [34] equips two-stream Transformer encoders [53] with multi-scale convolutions and achieves state-of-the-art performance in WiFi-based HAR. Video-based models. We apply 6 state-of-the-art video classification models for comparison, including ResNet [52], S3D [64], MViT-v1 [13], MViT-v2 [36], Swin-T [38], and Swin-S [38]. These models have demonstrated state-of-the-art performance in generic video classification.
4.2 Evaluation Metrics
We employ accuracy to measure the recognition performance of multi-user sensing, following the previous works [73, 7, 34]. Note that each multi-user sample includes multiple identity/location/activity labels corresponding to different users, respectively. For example, Fig. 0(a) shows a sample involving three identities: User 1, User 5, and User 6. Therefore, we measure the accuracy of recognizing these labels in each sensing task, rather than calculating the accuracy of classifying each sample. To evaluate model complexity and time efficiency, we adopt the number of parameters, floating point operations (FLOPs), and recognition throughput as metrics, following the previous works [34, 67].
4.3 Implementation Details
[…]
Table 2: Recognition performance of WiFi-based models on WiMANS in terms of accuracy (%). Models show desirable performance for multi-user identification and localization, while there is still vast room for improvement in activity recognition.
@@ 4.3 Implementation Details
[…]
ST-RF [73] 66.3±$\pm{1.33}$ 62.9±$\pm{0.36}$ 56.8±$\pm{0.14}$ 79.0±$\pm{1.32}$ 65.7±$\pm{0.55}$ 56.8±$\pm{0.09}$ 71.0±$\pm{2.20}$ 61.5±$\pm{0.54}$ 56.7±$\pm{0.09}$
MLP [67] 65.7±$\pm{0.91}$ 60.5±$\pm{0.47}$ 57.6±$\pm{0.13}$ 75.9±$\pm{0.62}$ 64.2±$\pm{0.47}$ 57.8±$\pm{0.16}$ 75.3±$\pm{0.59}$ 58.8±$\pm{0.26}$ 57.0±$\pm{0.12}$
LSTM [73] 80.7±$\pm{1.43}$ 66.9±$\pm{1.09}$ 59.2±$\pm{0.42}$ 89.5±$\pm{1.11}$ 71.7±$\pm{0.88}$ 59.0±$\pm{0.33}$ 86.7±$\pm{0.99}$ 67.8±$\pm{0.65}$ 57.7±$\pm{0.54}$
CNN-1D [58] 84.8±$\pm{0.80}$ 73.2±$\pm{0.22}$ 59.6±$\pm{0.24}$ 92.9±$\pm{0.26}$ 77.2±$\pm{0.19}$ 58.5±$\pm{0.11}$ 88.1±$\pm{0.46}$ 74.5±$\pm{0.31}$ 58.0±$\pm{0.17}$
CNN-2D [42] 90.0±$\pm{0.66}$ 77.0±$\pm{0.37}$ 59.5±$\pm{0.15}$ 96.4±$\pm{0.81}$ 82.1±$\pm{0.40}$ 59.0±$\pm{0.12}$ 91.1±$\pm{0.82}$ 79.1±$\pm{0.52}$ 58.2±$\pm{0.26}$
CLSTM [41] 90.2±$\pm{0.70}$ 73.0±$\pm{0.68}$ 61.8±$\pm{0.55}$ 94.6±$\pm{0.49}$ 77.4±$\pm{0.75}$ 60.9±$\pm{0.39}$ 92.8±$\pm{0.27}$ 71.8±$\pm{0.95}$ 61.0±$\pm{0.40}$
ABLSTM [7] 88.8±$\pm{1.36}$ 76.6±$\pm{0.42}$ 61.6±$\pm{0.50}$ 94.8±$\pm{0.55}$ 77.7±$\pm{0.61}$ 60.9±$\pm{0.27}$ 90.9±$\pm{0.54}$ 76.0±$\pm{0.55}$ 59.7±$\pm{0.15}$
THAT [34] 90.7±$\pm{0.51}$ 78.8±$\pm{0.74}$ 61.0±$\pm{0.32}$ 94.8±$\pm{0.59}$ 82.2±$\pm{0.84}$ 60.1±$\pm{0.28}$ 93.9±$\pm{0.88}$ 80.8±$\pm{1.11}$ 59.7±$\pm{0.29}$
[…]
ST-RF [73] 89.4±$\pm{1.08}$ 62.7±$\pm{0.57}$ 57.3±$\pm{0.08}$ 95.1±$\pm{0.83}$ 67.4±$\pm{0.65}$ 57.6±$\pm{0.13}$ 81.9±$\pm{1.26}$ 61.9±$\pm{0.56}$ 57.2±$\pm{0.21}$
MLP [67] 98.6±$\pm{0.13}$ 72.7±$\pm{0.82}$ 58.6±$\pm{0.14}$ 99.5±$\pm{0.11}$ 79.1±$\pm{0.42}$ 57.4±$\pm{0.14}$ 95.1±$\pm{0.58}$ 74.5±$\pm{0.87}$ 59.2±$\pm{0.23}$
LSTM [73] 98.9±$\pm{0.16}$ 76.1±$\pm{0.68}$ 60.6±$\pm{0.28}$ 99.7±$\pm{0.17}$ 81.3±$\pm{0.80}$ 59.5±$\pm{0.23}$ 95.8±$\pm{0.50}$ 77.4±$\pm{0.37}$ 59.8±$\pm{0.36}$
CNN-1D [58] 99.5±$\pm{0.21}$ 82.4±$\pm{0.34}$ 60.6±$\pm{0.22}$ 99.7±$\pm{0.12}$ 87.1±$\pm{0.32}$ 59.0±$\pm{0.16}$ 95.9±$\pm{0.36}$ 84.0±$\pm{0.58}$ 60.2±$\pm{0.14}$
CNN-2D [42] 99.1±$\pm{0.21}$ 82.7±$\pm{0.41}$ 59.8±$\pm{0.30}$ 99.8±$\pm{0.11}$ 88.2±$\pm{0.37}$ 58.6±$\pm{0.29}$ 95.3±$\pm{0.36}$ 84.4±$\pm{0.40}$ 59.6±$\pm{0.28}$
CLSTM [41] 99.7±$\pm{0.19}$ 82.0±$\pm{0.50}$ 64.2±$\pm{0.55}$ 99.8±$\pm{0.08}$ 88.1±$\pm{0.55}$ 62.1±$\pm{0.36}$ 97.6±$\pm{0.25}$ 84.4±$\pm{0.64}$ 64.8±$\pm{0.38}$
ABLSTM [7] 99.6±$\pm{0.29}$ 82.5±$\pm{0.67}$ 61.4±$\pm{0.22}$ 99.8±$\pm{0.08}$ 87.4±$\pm{0.25}$ 60.4±$\pm{0.16}$ 96.7±$\pm{0.38}$ 83.8±$\pm{0.53}$ 61.3±$\pm{0.15}$
THAT [34] 99.2±$\pm{0.19}$ 83.1±$\pm{0.67}$ 61.8±$\pm{0.29}$ 99.9±$\pm{0.00}$ 88.1±$\pm{0.59}$ 61.2±$\pm{0.37}$ 97.1±$\pm{0.29}$ 83.7±$\pm{1.05}$ 62.1±$\pm{0.41}$
[…]
ST-RF [73] 77.9±$\pm{1.05}$ 62.5±$\pm{0.39}$ 57.3±$\pm{0.09}$ 87.7±$\pm{0.91}$ 66.7±$\pm{0.34}$ 57.4±$\pm{0.07}$ 76.0±$\pm{1.06}$ 62.1±$\pm{0.28}$ 57.2±$\pm{0.08}$
MLP [67] 74.9±$\pm{0.48}$ 66.0±$\pm{0.32}$ 57.9±$\pm{0.13}$ 85.9±$\pm{0.78}$ 69.2±$\pm{0.28}$ 57.6±$\pm{0.07}$ 80.1±$\pm{0.53}$ 66.2±$\pm{0.48}$ 58.4±$\pm{0.08}$
LSTM [73] 86.9±$\pm{0.90}$ 69.5±$\pm{0.79}$ 59.6±$\pm{0.24}$ 93.9±$\pm{0.60}$ 75.2±$\pm{0.46}$ 59.7±$\pm{0.31}$ 90.1±$\pm{0.95}$ 70.4±$\pm{0.27}$ 58.9±$\pm{0.25}$
CNN-1D [58] 92.6±$\pm{0.42}$ 78.3±$\pm{0.25}$ 60.7±$\pm{0.14}$ 96.6±$\pm{0.21}$ 82.8±$\pm{0.25}$ 59.5±$\pm{0.17}$ 93.5±$\pm{0.21}$ 78.5±$\pm{0.21}$ 59.7±$\pm{0.12}$
CNN-2D [42] 94.6±$\pm{0.50}$ 79.2±$\pm{0.40}$ 59.2±$\pm{0.29}$ 98.0±$\pm{0.35}$ 83.6±$\pm{0.36}$ 58.5±$\pm{0.28}$ 94.5±$\pm{0.41}$ 80.1±$\pm{0.51}$ 58.3±$\pm{0.21}$
CLSTM [41] 93.1±$\pm{0.46}$ 79.9±$\pm{0.48}$ 64.0±$\pm{0.42}$ 97.4±$\pm{0.28}$ 83.1±$\pm{0.33}$ 62.4±$\pm{0.28}$ 95.3±$\pm{0.38}$ 78.2±$\pm{0.60}$ 63.5±$\pm{0.36}$
ABLSTM [7] 92.4±$\pm{0.39}$ 79.2±$\pm{0.38}$ 62.1±$\pm{0.16}$ 97.8±$\pm{0.17}$ 82.8±$\pm{0.42}$ 60.9±$\pm{0.17}$ 95.0±$\pm{0.34}$ 78.9±$\pm{0.54}$ 61.4±$\pm{0.29}$
THAT [34] 94.3±$\pm{0.51}$ 80.4±$\pm{0.56}$ 61.8±$\pm{0.19}$ 97.7±$\pm{0.33}$ 83.3±$\pm{0.45}$ 61.2±$\pm{0.22}$ 95.8±$\pm{0.40}$ 80.0±$\pm{0.82}$ 62.0±$\pm{0.33}$
4.4 Results and Analysis
WiFi-based Multi-user Sensing. Table 4.3 presents the multi-user sensing performance of WiFi-based models. (1) Using the 2.4 GHz WiFi band, THAT and CNN-2D are the best performing models for identification, while CNN-2D outperforms other models in localization, and CLSTM achieves the best performance in HAR. (2) With the 5 GHz WiFi band, CLSTM and THAT show the highest accuracy in recognizing multiple identities, while THAT and CNN-2D demonstrate their superiority in localization, and CLSTM obtains better results than other models in HAR. (3) Exploiting both 2.4 and 5 GHz WiFi bands, CNN-2D and THAT have the best results in identifying and localizing multiple users, while CLSTM yields the highest accuracy for HAR. Generally, we can observe that the state-of-the-art models show better performance on identification (90.69$\sim$99.99%) than localization (78.78$\sim$88.20%), while having even lower performance on HAR (60.87$\sim$64.82%). This is because HAR is more fine-grained than the other two tasks, causing higher difficulties to learn representative features. In contrast, multi-user identification is less fine-grained than localization and thereby results in better performance. Comparing different WiFi bands, using 5 GHz results in higher accuracy than 2.4 GHz, since the 2.4 GHz WiFi band suffers from more noise than 5 GHz, as mentioned in Section 3.2. Meanwhile, simply using the samples of two WiFi bands together cannot produce models of better performance. These results inspire us to study dual-band augmented multi-user sensing in the future.
@@ 4.4 Results and Analysis
Numbers of Users. Fig. 5 analyzes the impacts of user numbers on the performance of THAT. As expected, sensing more users simultaneously results in lower recognition accuracy. (1) In the classroom, when the maximum user number increases from 1 to 5, the accuracy of identification, localization, and HAR decreases by 5.70%, 15.39%, and 25.74%, respectively. (2) In the meeting room, sensing up to 5 users simultaneously diminishes the accuracy of three tasks by 2.26%, 14.59%, and 25.78% compared to sensing a single user. (3) In the empty room, similar results are evident where the accuracy of three tasks declines by 4.19%, 15.36%, and 25.90% as the maximum user number raises from 1 to 5. We can observe that, compared to localization and identification, the performance of HAR is more sensitive to changes in the number of users, because HAR is of finer granularity, and increasing users leads to more occlusion and mutual interference. Meanwhile, in varying environments, the number of users exerts distinct impacts on the recognition accuracy, which is worthwhile to further study. These results illustrate the usefulness of WiMANS and highlight the challenges of sensing multiple users simultaneously based on WiFi CSI.
Table 3: Recognition performance of video-based models on WiMANS in terms of accuracy (%). Video-based models show better performance than WiFi-based models.
@@
ResNet [52] 99.0±$\pm{0.93}$ 95.9±$\pm{1.70}$ 72.3±$\pm{0.86}$ 99.2±$\pm{0.94}$ 95.6±$\pm{5.14}$ 73.6±$\pm{1.13}$ 96.1±$\pm{1.39}$ 96.2±$\pm{5.38}$ 76.5±$\pm{4.74}$
S3D [64] 99.9±$\pm{0.00}$ 99.5±$\pm{0.50}$ 94.0±$\pm{1.35}$ 99.9±$\pm{0.04}$ 99.9±$\pm{0.02}$ 98.0±$\pm{0.12}$ 99.7±$\pm{0.32}$ 99.9±$\pm{0.01}$ 98.5±$\pm{0.15}$
MViT-v1 [13] 99.9±$\pm{0.05}$ 99.3±$\pm{0.26}$ 93.7±$\pm{0.44}$ 99.7±$\pm{0.12}$ 99.6±$\pm{0.05}$ 90.7±$\pm{0.11}$ 99.7±$\pm{0.04}$ 99.8±$\pm{0.01}$ 94.0±$\pm{0.10}$
MViT-v2 [36] 99.8±$\pm{0.06}$ 99.7±$\pm{0.05}$ 95.1±$\pm{0.11}$ 99.8±$\pm{0.13}$ 99.8±$\pm{0.04}$ 92.3±$\pm{0.13}$ 99.8±$\pm{0.12}$ 99.9±$\pm{0.01}$ 95.2±$\pm{0.11}$
Swin-T [38] 99.8±$\pm{0.05}$ 99.6±$\pm{0.05}$ 96.6±$\pm{0.12}$ 99.9±$\pm{0.05}$ 99.7±$\pm{0.07}$ 89.1±$\pm{0.14}$ 99.9±$\pm{0.04}$ 99.8±$\pm{0.01}$ 94.1±$\pm{0.13}$
Swin-S [38] 99.9±$\pm{0.07}$ 99.6±$\pm{0.06}$ 95.8±$\pm{0.11}$ 99.5±$\pm{0.08}$ 99.4±$\pm{0.05}$ 90.5±$\pm{0.11}$ 99.8±$\pm{0.12}$ 99.8±$\pm{0.02}$ 96.1±$\pm{0.11}$
Table 4: Model complexity and time efficiency on WiMANS, where WiFi-based models show their superiority over video-based models. (“Recs/s”: recognitions per second.)
@@
Video-based Multi-user Sensing. We provide the performance of video-based models in Table 3 for comparison. Overall, video-based models achieve better performance than WiFi-based models in three sensing tasks. Such results are expected since video-based analysis has been well studied in computer vision, which also demonstrate the quality of WiMANS, where users have performed distinguishing activities. Therefore, we can further exploit the videos in WiMANS as a reference for unexplored tasks in the future. Despite the remarkable performance of video-based models, they have much higher complexity and lower efficiency than WiFi-based models, as discussed below. Model Complexity and Time Efficiency. Table 4 highlights the superiority of WiFi-based models compared with video-based models in terms of model complexity and time efficiency. Except for MLP which is rarely used in practice due to excessive parameters, the maximum number of parameters (5.391 M for CLSTM) in WiFi-based models is 1.55$\times$ less than the minimum number of parameters (8.342 M for S3D) in video-based models. Similarly, the maximum FLOPs (3.208 G for ABLSTM) in WiFi-based models is 5.51$\times$ lower than the minimum FLOPs (17.670 G for ResNet) in video-based models. Moreover, the throughput of THAT is 16.33$\sim$17.02$\times$ higher than the throughput of S3D. These results signify that WiFi-based human sensing can achieve satisfactory performance with higher cost-effectiveness than video-based solutions. We further discuss the training time and testing time of models in Appendix C.